OpenStudy (anonymous):

USE THE DISTRIBUTIVE PROPERTY TO SIMPLIFY EXPRESSION -7(8x - 3) + x A. x + 10 B. -15x + 4 C. -9X - 4 D. -55x + 21

OpenStudy (anonymous):

-7 * 8 = -56, so we get -56x. -7 * -3 = +21, so our expression becomes: -56x +21 + x = -55x + 21

Parth (parthkohli):

-56x + 21 + x = -55x + 21